# When GPUs Fail Quietly: Observability-Aware Early Warning Beyond Numeric Telemetry

### 17 Mar 2026 v1

@@ IV-A Telemetry from GWDG GPU Nodes
Slurm is configured to run the LBNL Node Health Check (NHC) (https://github.com/mej/nhc) every 30 min to check nodes for failures not caught by jobs crashing and getting stuck in the epilog with “Kill task failed”. The NHC script, which determines which tests are done, checks number number of CPU cores, amount of RAM, network interfaces, filesystem mounts, etc. as well as GPUs. The NHC script is how much GPU failures are caught. Any problems found by NHC result in Slurm draining the node which shows up as a node state transition which is collected by the Prometheus Slurm Exporter. The following GPU checks are done:
@@ IV-B Incident Catalog and Windowing
In addition to continuous telemetry, we use an operator-curated incident catalog to anchor analysis windows around operationally relevant events. The catalog entries originate from operational monitoring and scheduler-side evidence, including Slurm node-state transitions and system-level logs associated with GPU-related failures. The incident catalog therefore provides coarse failure annotations rather than precise machine-aligned labels.
Each incident entry specifies the affected node, incident date which is either when it happened or when it was noticed (e.g. failure on a Saturday not noticed till Monday), free-text description, failure category, and asymmetric collection bounds (beforeHours, afterHours), which define the raw telemetry interval (collectStart, collectEnd) used for analysis.
@@ V-A Windowing and Notation
Here, $\mathbf{x}^{gpu}_{n,g}(t)$ is GPU-local telemetry for GPU $g$ on node $n$, $\mathbf{x}^{pipe}_{n}(t)$ is node-level monitoring-pipeline (observability) features shared by all GPUs on node $n$, and $\mathbf{x}^{os}_{n}(t)$ denotes node-level OS telemetry.
@@
We distinguish three time scales used throughout the analysis: (i) incident collection windows defined by the operator catalog (beforeHours/afterHours, e.g., 24 h/2 h), which determine the raw telemetry interval extracted around each incident; (ii) detector feature windows with length $w$ and stride $s$, used to aggregate time series for anomaly detection; and (iii) a compact forensic comparison window consisting of a $30\text{\,}\mathrm{m}\mathrm{i}\mathrm{n}$ baseline interval and a $5\text{\,}\mathrm{m}\mathrm{i}\mathrm{n}$ interval adjacent to $t_{0}$ (as configured by tAfterMin), used to rank metric shifts and structural disappearance around detachment events. This separation avoids conflating data collection bounds with detector aggregation and forensic comparison.
@@ V-D Monitoring Pipeline Feature Plane
We quantify monitoring pipeline behavior using scrape- and probe-level indicators when available in the selected archive subset(s). Unlike conventional monitoring practice, where missing or incomplete telemetry is often treated purely as a data quality issue, we explicitly model monitoring degradation as a first-class anomaly signal.
@@ VI-B Weak-Event Construction for Unlabeled Evaluation
This proxy is intended to capture drift-dominated instability regimes with gradual numeric deviation before operational impact. 8It is not intended to represent detachment-class failures, where the dominant pre-fault manifestation is structural observability degradation rather than a stable excursion in conventional GPU telemetry.
@@ VI-C Empirical Weak-Event Characteristics
To quantify pre-fault feature behavior in the weak-event slice, we summarize incident-anchored feature shifts extracted from weak_events_summary.csv. Each row corresponds to a pre-failure window aligned to an operator incident time $t_{0}$. The table reports the number of available feature signals (numSignalsLong) and the dominant feature shifts ranked by absolute delta within the forensic comparison window.
@@
Across the weak-event slice, variance-shift statistics are frequently zero and do not provide stable ranking. Observable pre-fault differences are therefore dominated by delta-level feature shifts and structural feature disappearance rather than sustained numeric drift. This empirical property of the slice explains why drift-based weak-event lead time is limited in the baseline results.
@@ VI-D Incident-Anchored Evaluation for Detachment Failures
Because detachment-class failures are not captured by signature-derived weak events, we additionally perform incident-anchored evaluation using the operator incident catalog. For these cases, the reference time $t_{0}$ is aligned via scrape payload collapse (scrapeCountDrop). Detectors are evaluated on whether observability-aware indicators trigger within the incident window before operational impact is recorded.
@@ VII Results and Case Studies
We report detachment-class case studies from nodes ggpu142 and ggpu149, spanning February 2025 to January 2026, as identified in the incident catalog. Notably, for ggpu149 on 12 June 2025, the operator incident record indicates that NHC detected the condition many hours after the underlying GPU detachment event. Forensic alignment shows that GPU telemetry disappearance and scrape degradation occurred near $t_{0}$, well before operator detection. This gap illustrates the operational importance of observability-aware indicators that react to telemetry collapse earlier than delayed health checks.
@@ VII-A Baseline Results Under Fixed Alert Budget
Table VI shows that using both GPU and observability features together (“Joint”) gives earlier warnings than using GPU data alone. In particular, Joint Isolation Forest achieves the highest average lead time (7 windows) and the longest maximum lead (29 windows). GPU-only detectors typically trigger close to or after event onset (median lead = 0), meaning they often detect instability late. Having this, combining observability signals with GPU telemetry improves early detection without increasing the alert budget.
@@ VIII-B Dataset Choice and External Validity
Our choice of using GWDG production telemetry rather than public datasets such as Marconi100 ExaData is driven by the need for cross-layer correlation between telemetry and scheduler-level node state. While ExaData offers valuable large-scale GPU measurements, it does not expose Slurm (or equivalent) logs indicating node drain, node down, or failure-reason events that can be temporally aligned with GPU, power, and node-level metrics. This limits the ability to assess whether detected anomalies precede operationally meaningful node state transitions.
@@ VIII-E Recurrence and Host-Level Hazard
From a reliability engineering perspective, recurrence is a more informative hazard signal than the severity of any single event. Nodes that repeatedly experience device detachments are unlikely to self-heal and pose an elevated risk to workload stability and operational efficiency. Consequently, anomaly detection systems for sensitive infrastructures should incorporate host-history aggregation and recurrence-aware scoring, enabling proactive intervention such as node quarantine, hardware replacement, or retirement before catastrophic failure occurs.

### 04 Apr 2026 v2

@@ IV-A Telemetry from GWDG GPU Nodes
Slurm is configured to run the LBNL Node Health Check (NHC) (https://github.com/mej/nhc) every $30\text{\,}\mathrm{m}\mathrm{i}\mathrm{n}$ to check nodes for failures not caught by jobs crashing and getting stuck in the epilog with “Kill task failed”. The NHC script, which determines which tests are done, checks the number of CPU cores, amount of RAM, network interfaces, filesystem mounts, etc. as well as GPUs. The NHC script is how most GPU failures are caught. Any problems found by NHC result in Slurm draining the node which shows up as a node state transition which is collected by the Prometheus Slurm Exporter. The following GPU checks are done:
@@ IV-B Incident Catalog and Windowing
In addition to continuous telemetry, we use an operator-curated incident catalog to anchor analysis windows around operationally relevant events. The failures in the catalog originate from real operational monitoring and scheduler-side evidence, including Slurm node-state transitions (node health script checks GPUs) and system-level logs associated with GPU-related failures. The incident catalog therefore provides coarse failure annotations rather than precise machine-aligned labels.
Each incident entry specifies the affected node, incident date which is either when it happened or when it was noticed (e.g. failure on a Saturday not noticed till Monday), free-text description, failure category, and asymmetric collection bounds (beforeHours, afterHours), which define the raw telemetry interval (collectStart, collectEnd) used for analysis. Since the incident date in the catalog could be on the day of the error or when it was noticed, pre-processing had to be done to find when the incident took place more precisely to determine the collectStart and collectEnd to use. For each incident entry in the catalogue,
  1. 1.
    The times of all slurm node-state transitions on the from OK (idle, alloc, mix) to failure (drain, draining, down, no response, rebooting) were found. The failures either come from the node health script check which checks the GPUs empirically (every $30\text{\,}\mathrm{m}\mathrm{i}\mathrm{n}$), a GPU job failing, or the node no longer being responsive.
  2. 2.
    If there is at least one transition from OK to failure on the day logged in the incident catalog, the first one is chosen as the incident time.
  3. 3.
    If there is at least one transition from OK to failure in the 3 days prior to the day logged in the incident catalog, the last one is chosen as the incident time.
  4. 4.
    Otherwise the incident is discarded due to not being able to identify when the incident occurred (most likely this is from not logging the incident in the catalog till more than 3 days after it occurred).
@@ V-A Windowing and Notation
Here, $\mathbf{x}^{gpu}_{n,g}(t)$ is GPU-local telemetry for GPU $g$ on node $n$, $\mathbf{x}^{pipe}_{n}(t)$ is node-level monitoring-pipeline (observability) features shared by all GPUs on node $n$, and $\mathbf{x}^{os}_{n}(t)$ denotes node-level OS telemetry. We define $t_{0}$ as the incident time.
@@
We distinguish three time scales used throughout the analysis: (i) incident collection windows defined by the operator catalog after the proprocessing stop to more precisely find when the incident occurred (beforeHours/afterHours, e.g., 24 h/2 h), which determine the raw telemetry interval extracted around each incident; (ii) detector feature windows with length $w$ and stride $s$, used to aggregate time series for anomaly detection; and (iii) a compact forensic comparison window consisting of a $30\text{\,}\mathrm{m}\mathrm{i}\mathrm{n}$ baseline interval and a $5\text{\,}\mathrm{m}\mathrm{i}\mathrm{n}$ interval adjacent to $t_{0}$ (as configured by tAfterMin), used to rank metric shifts and structural disappearance around detachment events. This separation avoids conflating data collection bounds with detector aggregation and forensic comparison.
@@ V-D Monitoring Pipeline Feature Plane
We quantify monitoring pipeline behavior using scrape- and probe-level indicators when available in the selected archive subset(s). Unlike conventional monitoring practice where missing or incomplete telemetry is often treated purely as a data quality issue, we explicitly model monitoring degradation as a first-class anomaly signal.
@@ VI-B Weak-Event Construction for Unlabeled Evaluation
This proxy is intended to capture drift-dominated instability regimes with gradual numeric deviation before operational impact. It is not intended to represent detachment-class failures, where the dominant pre-fault manifestation is structural observability degradation rather than a stable excursion in conventional GPU telemetry.
@@ VI-C Empirical Weak-Event Characteristics
To quantify pre-fault feature behavior in the weak-event slice, we summarize incident-anchored feature shifts extracted from weak_events_summary.csv. Each row corresponds to a pre-failure window aligned to $t_{0}$, the operator incident time after the pre-processing search for the actual failure time. Table III reports the number of available feature signals (numSignalsLong) and the dominant feature shifts ranked by absolute delta within the forensic comparison window.
@@
Across the weak-event slice, variance-shift statistics are frequently zero and do not provide stable ranking. Observable pre-fault differences are therefore dominated by delta-level feature shifts and structural feature disappearance rather than sustained numeric drift. This empirical property of the slice explains why drift-based weak-event lead time is limited in the baseline results. Additionally, the dominant delta-level features are often just changes in the host CPU load or RAM usage, rather than GPU metrics. These are potentially from a job allocating a lot of memory right before failure or memory deallocation right around the failure, possibly due to the job failing due to the GPU failure and the processes dying.
@@ VI-D Incident-Anchored Evaluation for Detachment Failures
Because detachment-class failures are not captured by signature-derived weak events, we additionally perform incident-anchored evaluation using the operator incident catalog after pre-processing to find the approximate incident time. For these cases, the reference time $t_{0}$ is aligned via scrape payload collapse (scrapeCountDrop). Detectors are evaluated on whether observability-aware indicators trigger within the incident window before operational impact is recorded.
@@ VII Results and Case Studies
We report detachment-class case studies from nodes ggpu142 and ggpu149, spanning February 2025 to January 2026, as identified in the incident catalog in Table V. Notably, for ggpu149 on 12 June 2025, the operator incident record indicates that NHC detected the condition many hours after the underlying GPU detachment event. Forensic alignment shows that GPU telemetry disappearance and scrape degradation occurred near $t_{0}$, well before operator detection. This gap illustrates the operational importance of observability-aware indicators that react to telemetry collapse earlier than delayed health checks.
@@ VII-A Baseline Results Under Fixed Alert Budget
Table VI shows that using both GPU and observability features together (“Joint”) gives earlier warnings than using GPU data alone. In particular, Joint Isolation Forest achieves the highest average lead time (7 windows) and the longest maximum lead (29 windows). GPU-only detectors typically trigger close to or after event onset (median lead = 0), meaning they often detect instability late. Having this, combining observability signals with GPU telemetry improves early detection without increasing the alert budget. The lead time for joint IF and OCSVM are potentially sufficient that if they could be extended to a live monitoring system, compute jobs on a node right before GPU failure could be signaled to warn them that GPU failure is imminent so that suitably designed jobs have a chance to take snapshots of their current progress which they could later resume in another job on a different node, potentially even submitting the follow-up job right from the job as it is cleaning up (not all workloads support snapshots, but many do). Though getting users to design their jobs to support this might actually be the harder part.
@@ VIII-B Dataset Choice and External Validity
Our choice of using GWDG production telemetry rather than public datasets such as Marconi100 ExaData is driven by the need for cross-layer correlation between telemetry and scheduler-level node state derived from compute job failures and empirically checking the GPUs by active probing (nvidia-smi and allocating VRAM). While ExaData offers valuable large-scale GPU measurements, it does not expose Slurm (or equivalent) logs indicating node drain, node down, or failure-reason events that can be temporally aligned with GPU, power, and node-level metrics. This limits the ability to assess whether detected anomalies precede operationally meaningful node state transitions.
@@ VIII-E Recurrence and Host-Level Hazard
From a reliability engineering perspective, recurrence is a more informative hazard signal than the severity of any single event. Nodes that repeatedly experience device detachments are unlikely to self-heal and pose an elevated risk to workload stability and operational efficiency. Consequently, anomaly detection systems for sensitive infrastructures should incorporate host-history aggregation and recurrence-aware scoring, enabling proactive intervention such as node quarantine, reallocating nodes to lower priority and shorter duration workloads that are easier to redo (e.g. short runs for testing codes before longer computations), hardware derating (e.g. reducing the maximum clock frequencies of some components so that they can still operate but at reduced performance if that reduces the chance of failure), hardware replacement, or retirement before catastrophic failure occurs.